OpenStudy (brooke..help00):

The equation below shows the relationship between the temperature in degrees Celsius, C, and degrees Fahrenheit, F: C = 5 over 9(F − 32) Which of the following formulas correctly solves for F? F = 9 over 5C + 32 F = 9 over 5C − 32 F = 9C + 32 over 5 F = 9C − 32 over 5

OpenStudy (brooke..help00):

@zepdrix

OpenStudy (mathmale):

I'd suggest that you actually solve the given equation for F. To eliminate fractions, mult. both sides of the given eq'n by 9. Show your results.

OpenStudy (brooke..help00):

That's the part I'm having trouble with...

OpenStudy (mathmale):

surely you can multiply that C by 9. Try it. Your result?

OpenStudy (brooke..help00):

I stink at math :/ This I so confusing... :(

OpenStudy (mathmale):

Such statement do not help you and do not help me. You have seen this math before and should have little or no trouble with it.

OpenStudy (mathmale):

The given equation begins with C. Multiply that by 9.

OpenStudy (brooke..help00):

\[\frac{ 9 }{ 5 } \div C=?\]

OpenStudy (brooke..help00):

I meant the multiply sign

OpenStudy (mathmale):

I gave you C and asked that you multiply it by 9. Try again.

OpenStudy (mathmale):

If I give you 5 and ask you to mult. that by 3, you'd write 3(5) = 15. If I give you C and ask you to mult. that by 9, you'd write ... what?

OpenStudy (brooke..help00):

3(C)

OpenStudy (mathmale):

but wouldn't 9 times c be 9c? Why the 3?

OpenStudy (brooke..help00):

Oh sorry I thought I put 9

OpenStudy (mathmale):

C = (5 over 9)(F − 32) You have mult. that C by 9 and gotten 9C. Now multiply that (5 over 9) by 9. Your result? (You have to keep that (F-32) as part of the resulting equation.)

OpenStudy (mathmale):

Left side: 9C Right side: 9(5/9)(F-32) Please simplify this.

OpenStudy (mathmale):

Please simplify:\[9(\frac{ 5 }{ 9 })(F-32)\]

OpenStudy (brooke..help00):

Would it be 5?

OpenStudy (brooke..help00):

5C

OpenStudy (mathmale):

The 9 cancels out, yes. That leaves you with 5(F-32). You cannot drop or forget the (F-32). You began with C = 5 over 9(F − 32) and now you have 9c=5(F-32). Divide both sides by 5, so that you'll have (F-32) on the right side. please show your work.

OpenStudy (mathmale):

Left side: divide 9C by 5. Type out your result.

OpenStudy (brooke..help00):



OpenStudy (mathmale):



OpenStudy (mathmale):

I need your response, very soon, if y ou want me to continue here.

OpenStudy (brooke..help00):

You would divide and it would become 6.4 right>?

OpenStudy (mathmale):

Divide 9C by 5. All you have to do is retype this into one fraction.

OpenStudy (mathmale):

Type 9C in the numerator of your fraction and 5 in the denom.

OpenStudy (brooke..help00):

9/5C

OpenStudy (mathmale):

9C/5 would be better. Now divide 5(F-32) by 5.

OpenStudy (brooke..help00):

6.4?

OpenStudy (mathmale):

You have [ 5 times (F-32) ] and need to divide this whole expression by 5. Try again. Do NOT change or drop the (F-32) part.

OpenStudy (brooke..help00):

What?

OpenStudy (mathmale):

I can't type that any clearer than I already have. Think about this and then answer my question.

OpenStudy (mathmale):

Simplify \[\frac{ 5(F-32) }{ 5 }\]

OpenStudy (mathmale):

As I said before, leave the (F-32) part alone. Focus on the 5/5. Your result?

OpenStudy (brooke..help00):

Just the 5/5 which would cancel it out. You have F-32 keft?

OpenStudy (mathmale):

yes.

OpenStudy (mathmale):

so now you have\[\frac{ 9C }{ 5 }=F-32\]

OpenStudy (mathmale):

Your goal is to solve for F. You must isolate F on the right side of your equation. How would you do this? Solve for F, please.

OpenStudy (brooke..help00):

add 32

OpenStudy (mathmale):

add 32 to both sides. Do this now, please.

OpenStudy (brooke..help00):

How would you add it to your fraction???

OpenStudy (mathmale):

Write 9C/5 and then write "+32" after it. Write out your whole new equation now, please.

OpenStudy (brooke..help00):

9C/5 + 32 = F

OpenStudy (mathmale):

Right. That's your formula. Does it match any of your answer choices?

OpenStudy (mathmale):

Brooke?

OpenStudy (lowkey.s):

She's offline